OpenStudy (anonymous):

Which number equals 255

OpenStudy (anonymous):

\[\sqrt{255}\]

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \sqrt{5 *3 * 17} }\)

OpenStudy (anonymous):

?............... \[10\sqrt{2}\] \[15\sqrt{5} \] 15 25 this is the choices

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Wait

Parth (parthkohli):

Closest is 15. Although not exactly.

OpenStudy (anonymous):

thank you!